OpenStudy (1dfanatic):

Use the solution of the system below to find x - y. (So, solve the system, then subtract your values you got.) 4x - 2y = 11 3x - 4y = -6

jhonyy9 (jhonyy9):

do you wan to solve it by elimination or by substitution - what method ?

OpenStudy (1dfanatic):

Um substitution i guess

jhonyy9 (jhonyy9):

ok how many will be y from the first equation ?

jhonyy9 (jhonyy9):

in case of the first equation what will get for y= ?

OpenStudy (1dfanatic):

y=2x-11/2..?

jhonyy9 (jhonyy9):

right so than by substitution method you need substituting this value of y in place of y inside second equation - yes ?

jhonyy9 (jhonyy9):

and so will get the value of x

OpenStudy (1dfanatic):

x=2(2y-3)/3

jhonyy9 (jhonyy9):

no you need getting value without y

jhonyy9 (jhonyy9):

4x - 2y = 11 => y=(4x-11)/2 - yes ? 3x - 4y = -6

jhonyy9 (jhonyy9):

so this value of y you need substituting in place of y inside second equation - what will get ?

jhonyy9 (jhonyy9):

3x -4(4x-11)/2 = -6 do you can continue it ?

OpenStudy (1dfanatic):

x=28/5

jhonyy9 (jhonyy9):

the 4 from the front of parentheses and the denominator simplifie it by 2 and will get

jhonyy9 (jhonyy9):

so ok. than this is right value of x so just substitute it in place of x and calcule the value of y - ok ?

jhonyy9 (jhonyy9):

x=28/5 y=(4x-11)/2 so y = (4*28/5 -11)/2 y= ?

OpenStudy (1dfanatic):

57/10?

jhonyy9 (jhonyy9):

so than this is right so your exercise aboce said that at the end when you have got the value of x and y you need to find x-y= ?

OpenStudy (1dfanatic):

-0.1

jhonyy9 (jhonyy9):

x-y= ? x= 28/5 y= 57/10 x-y= 28/5 - 57/10 = ? yes -0,1

jhonyy9 (jhonyy9):

wsa my pleasure bye bye